OpenStudy (anonymous):

what is -3 &4/7 divided by 1/5

OpenStudy (anonymous):

-15 and 20/7

OpenStudy (anonymous):

\[-3\frac{4}{7} * \frac{1}{5}\] \[=-\frac{25}{7}*\frac{1}{5}\] \[=-\frac{5}{7}\]